# From Optimization to Prediction: Transformer-Based Path-Flow Estimation to the Traffic Assignment Problem

### v1

@@ Abstract
The traffic assignment problem is essential for traffic flow analysis, traditionally solved using mathematical programs under the Equilibrium principle. These methods become computationally prohibitive for large-scale networks due to non-linear growth in complexity with the number of OD pairs. This study introduces a novel data-driven approach using deep neural networks, specifically leveraging the Transformer architecture, to predict equilibrium path flows directly. By focusing on path-level traffic distribution, the proposed model captures intricate correlations between OD pairs, offering a more detailed and flexible analysis compared to traditional link-level approaches. The Transformer-based model drastically reduces computation time, while adapting to changes in demand and network structure without the need for recalculation. Numerical experiments are conducted on the Manhattan-like synthetic network, the Sioux Falls network, and the Eastern-Massachusetts network. The results demonstrate that the proposed model is orders of magnitude faster than conventional optimization. It efficiently estimates path-level traffic flows in multi-class networks, reducing computational costs and improving prediction accuracy by capturing detailed trip and flow information. The model also adapts flexibly to varying demand and network conditions, supporting traffic management and enabling rapid ‘what-if’ analyses for enhanced transportation planning and policy-making.
@@ 1 Introduction
The Traffic Assignment Problem (TAP) is a process of determining the propagation of flows over the transportation network. The goal is to calculate the network state, given the travel demand between various origin-destination (OD) pairs and the network’s capacity constraints (Yosef Sheffi. Prentice-Hall, 1986). Traditionally, this problem is solved through mathematical programs under the User Equilibrium (UE) principle, which assumes drivers possess perfect information and make fully rational choices (Wardrop, 1952). Despite potential deviations from reality, this approach consistently provides reasonable solutions to the traffic assignment problem (Bar-Gera, 2002; Jafari et al., 2017). However, the computation for determining optimal solutions in large traffic networks is prohibitively costly. This is because the problem’s complexity grows non-linearly with the increase in the number of OD pairs and directly depends on feasible paths. When the size of the network (the number of links and nodes in a representative graph) increases, allowing us to explore more paths, the number of feasible paths also increases, and the OD demand matrix may grow accordingly, leading to a non-linear increase in computation time (Patriksson, 2015). Because changing the input (demand and graph) results in a new problem to be solved, it requires re-calculation. Therefore, there is a critical need to find another way to capture the complex nature of traffic flow and demand in large networks in order to use the already-solved problem to estimate the solution to the new problem.
@@
Mathematically, the UE, is typically calculated using optimization methods such as the gradient method or fixed-point algorithm (Tan et al., 2020; Mehrabipour and Hajbabaie, 2021; Liu et al., 2023), e.g., Frank-Wolfe algorithm Fukushima (1984). However, with the presence of data-driven approaches and the application of machine learning, an alternative solution to this problem could be prediction instead of optimization. In other words, we can employ supervised learning to learn the optimal solution and predict it given a new demand or a new graph. Recently, a few studies have tried to estimate the link flow solution of UE using machine learning (W. et al., 2023; Rahman and Hasan, 2023; Liu and Meidani, 2024; Liu et al., 2025; Liu and Meidani, 2025; Hu and Xie, 2025) with graph neural networks (GNNs) emerging as a particularly popular approach for utilising network topology to capture spatial dependencies. For example, Rahman and Hasan (2023); Liu et al. (2025) uses graph convolutional neural networks (GCN) to forecast traffic flow patterns, while Liu and Meidani (2024); Hu and Xie (2025) employed graph attention networks (GATs) to accelerate assignment across large-scale planning scenarios. While promising, GNNs are constrained by their reliance on local message passing, which limits their ability to model long-range dependencies. Deeper stacking can address this, but at the cost of higher computation and over-smoothing of features.
@@
Other approaches have pursued surrogate formulations of the traffic assignment problem. For instance, Zhang and Qian (2025) introduced a low-rank approximation for path-based network models, generating hypothetical path sets based on flows rather than topological structure. Liu et al. (2023) proposes an end-to-end learning framework that embeds the traffic equilibrium problem as a Variational Inequality (VI) layer, producing flows through an iterative equilibrium procedure. Yet these methods remain tied to surrogate calibration or traditional iterative equilibrium procedures.
@@
To address these limitations, we introduce a Transformer-based framework for TAP that employs global self-attention, enabling each element of the input sequence to dynamically weigh its relevance to all others and capture complex, long-range dependencies across the network. The attention mechanism allows the model to selectively prioritize OD pairs with nonzero demand or links with critical features while maintaining global context, enhancing learning efficiency. In contrast, prior TAP models using GAT-based encoder-decoder architectures restrict attention to local neighborhoods and decode flows through MLP layers, limiting their ability to model global interactions (Hu and Xie, 2025).
Most existing approaches emphasize link-level predictions, leaving path-level demand propagation underexplored despite its importance for understanding traveler behavior. The proposed model predicts detailed path-level flows, enabling an understanding of demand propagation and facilitating flexible ‘what-if’ analyses, i.e., scenario-based evaluations of the effects of simulated changes in network topology or OD demand. By predicting flows at the path level, the model captures the interactions between different OD pairs, overcoming the limitations of link-based approaches that fail to account for these interdependencies. The model learns this mapping in a fully data-driven manner by observing feasible path flows that inherently respect flow conservation and link capacity constraints, producing physically and operationally feasible network states.
Existing neural network approaches are mostly trained and evaluated on familiar scenarios and in single-class networks, leaving their performance under new or unforeseen traffic or network conditions largely unexplored. While Deep Neural Networks (DNNs) have proven effective in domains such as computer vision and NLP Wang et al. (2022), and are increasingly applied to transportation forecasting, the use of sequence-based architectures, particularly Transformers, remains largely unexplored in TAP (Wen et al., 2023; Xue et al., 2025; Jia et al., 2025). This study addresses these gaps by estimating equilibrium flows in multi-class networks and across diverse demand patterns and network-change scenarios, offering both efficiency and robustness in complex traffic scenarios. To the best of our knowledge, this is the first TAP framework to employ a Transformer-based sequence learning approach incorporating both encoder and decoder global attention.
@@
To summerize, the major contributions of this work are as follows: (i) A Transformer-based model is proposed for directly learning and predicting UE path flow distributions in both single- and multi-class traffic networks; (ii) The model incorporates both network topology and detailed OD demand information, allowing it to generalize to unseen demand patterns and network changes; (iii) Our approach eliminates the computationally intensive optimization step, providing solutions orders of magnitude faster than traditional methods; and (iv) The model supports robust ‘what-if’ scenario analysis under extreme perturbations, where link-removal based topology changes, missing-link observation cases, as well as multi-class and large-scale network applications particularly in terms of OD demand are considered jointly with incomplete OD demand.
@@ 2 Related literature
Beginning with Wardrop’s UE principle and the convex optimisation formulation of Beckmann et al, these approaches laid the foundation of the TAP (Wardrop, 1952; Beckman and McGuire, 1956). To reduce the need for full path enumeration, some analytical approaches developed link-based algorithms, which directly solve for link flows (Fukushima, 1984), as well as column and bush-based methods, which aim to find only promising subsets of paths to improve efficiency (Ford and Fulkerson, 1958; Bar-Gera, 2002; Dial, 2006). Despite these advances, TAP remains computationally intensive for large networks, where evaluating thousands of ‘what-if’ scenarios can take days. This motivates the shift toward supervised machine learning models trained on solver-generated solutions, offering significantly faster inference while retaining the accuracy of the underlying optimization framework.
@@ 2.2 Deep Learning (DL) models
Besides the calculation of UE by optimization approaches, there are multiple deep learning approaches for traffic estimation, which can be related to estimating the state of the network. RNNs have been widely used to model temporal dependencies in traffic flow data for prediction. Studies like those by Bruno et al. (2020) and Wu et al. (2023) have shown the ability of Long Short Term Memory (LSTM) networks, a type of RNN, to capture long-term and short-term dependencies in traffic patterns. However RNNs are recursive and cannot parallelize all sequences in parallel, and they are weak for variable-length data (Ma et al., 2015; Bogaerts et al., 2020). W. et al. (2023) had built a CNN-based model to estimate link flow from incomplete OD data. However, this model was applied on a small-sized dataset of 1000 data points and it does not address the estimation tasks of path flow distribution.
@@ 3 Comparison Between the Proposed Model and Other Neural Network Models
Model Type
Model Output
Baseline
OD
Data
Perturbations Conservation Handling
@@
As discussed in previous sections, the application of neural networks to TAP remains relatively underexplored. Table 1 compares our approach with other state-of-the-art neural network based traffic assignment models in terms of model outputs, handling of perturbations, and flow conservation strategies.
@@
While link or path-level flow conservation ensures locally feasible flows, it does not necessarily guarantee that OD demand is satisfied. To address this, Hu and Xie (2025) applied a post-processing step to adjust path flows and enforce OD demand consistency. In contrast, our path-based Transformer model, trained on ground-truth path flows generated under equilibrium conditions, predicts flows that inherently sum to each OD pair’s demand, ensuring OD-level conservation without any additional adjustment. Because OD conservation is implicit and learned, no iterative enforcement or adjustment is needed at inference, making predictions potentially faster.
@@ 4.1 Problem formulation
As mentioned before, the goal is to estimate the optimal solution of path flow distribution from given OD demands and the graph instead of solving it. In mathematical terms, we consider a network represented as a graph $G=(V,E)$, where $V$ stands for the set of nodes with size $N$, and $E$ for the set of links. Let $R$ represent the set of OD pairs. For each OD pair, denoted by $r\in R$, where $r=(v_{1},v_{2}),\forall v_{1},v_{2}\in V$, vehicles of class $z\in Z$ can travel via several finite and nonempty feasible paths of path set $\mathbf{P}_{r}$. Let $n$ denote the number of classes in the network. Each set $\mathbf{P}_{r}\in\mathbf{P}^{k}$ encompasses $k$ feasible paths for each OD pair $r$. Furthermore, let $\mathbf{X}$ represents OD demand matrix, $x^{z}_{r}\in\mathbf{X}$ represents the demand of OD pair $r$ traveled by class $z$, and $f^{r,z}_{\mathbf{P}_{r}}\in\mathbf{F}$ represents paths flow distribution of OD pair $r$ on path set $\mathbf{P}_{r}$, with class $z$. Each set $f^{r,z}_{\mathbf{P}_{r}}$ encompasses $k$ flow values for every OD pair $r$.
@@
The ultimate goal is to find $\mathbf{F}^{*}$, the optimal path flow distribution for UE condition defined by Wardrop (1952). Each set of path flow distribution ${f^{r,z}_{\mathbf{P}^{r}}}^{*}\in\mathbf{F}^{*},\forall r,\forall z$ is the optimal solution for OD pair $r$ on path set $p_{r}$ with class $z$. Equation 1 shows that for all classes $z\in Z$, the paths connecting any OD pair $r$ can be categorized into two groups: those that carry flow, where the travel time $c^{r,z}_{p}$ matches the minimum OD travel time $u^{z}_{r}$, and those that do not carry flow, where the travel time is greater than or equal to the minimum OD travel time. Equation 2 ensures that $u^{z}_{r}$ is the minumum path cost. Equation 3 ensures that all the demand is served. Equation 4 ensures the positivity of variables. However, this UE solution is only applicable when all drivers have perfect knowledge of the network state. In reality, it is not the case, so normally there will be a delay in the network. In the literature, this imperfection of knowledge is represented by stochastic terms that lead to other equilibrium terms such as Stochastic User Equilibrium (SUE) or Boundedly Rational UE (BRUE) (Ameli et al., 2020).
@@ 5 Numerical Experiments
Two numerical experiments are conducted to evaluate the accuracy, efficiency, and generalization capability of the proposed framework. In Section 5.3, We first experiment with a synthetic Manhattan-like network to assess the basic effectiveness of the model. We evaluate the model’s performance under conditions of incomplete OD demand. The second experiment in Section 5.4 is on urban transportation networks (Sioux Falls and Eastern-Massachusetts network) with several ‘what-if’ scenarios to examine robustness against various realistic network perturbation scenarios. The implementation details and datasets used in these experiments are available at https://github.com/VivianeLe/Path_Flow_Prediction. The details of the experiments will be explained in the following sections.

### v2

@@ Abstract
The traffic assignment problem is essential for traffic flow analysis, traditionally solved using mathematical programs under the Equilibrium principle. These methods become computationally prohibitive for large-scale networks due to non-linear growth in complexity with the number of OD pairs. This study introduces a novel data-driven approach using deep neural networks, specifically leveraging the Transformer architecture, to predict equilibrium path flows directly. By focusing on path-level traffic distribution, the proposed model captures intricate correlations between OD pairs, offering a more detailed and flexible analysis compared to traditional link-level approaches. The Transformer-based model drastically reduces computation time, while adapting to changes in demand and network structure without the need for recalculation. Numerical experiments conducted on the Manhattan-like synthetic network, the Sioux Falls network, the Eastern-Massachusetts network, and the Anaheim network demonstrate that the proposed model is orders of magnitude faster than conventional optimization. It efficiently estimates path-level traffic flows in multi-class networks, reducing computational costs and improving prediction accuracy by capturing detailed trip and flow information. The model also adapts flexibly to varying demand and network topology conditions, supporting traffic management and enabling rapid ‘what-if’ analyses for enhanced transportation planning and policy-making.
@@ 1 Introduction
The Traffic Assignment Problem (TAP) is a process of determining the propagation of flows over the transportation network. The goal is to calculate the network state, given the travel demand between various origin-destination (OD) pairs and the network’s capacity constraints (Sheffi, 1985). Traditionally, this problem is solved through mathematical programs under the User Equilibrium (UE) principle, which assumes drivers possess perfect information and make fully rational choices (Wardrop, 1952). Despite potential deviations from reality, this approach consistently provides reasonable solutions to the traffic assignment problem (Bar-Gera, 2002; Jafari et al., 2017). However, the computation for determining optimal solutions in large traffic networks is prohibitively costly. This is because the problem’s complexity grows non-linearly with the increase in the number of OD pairs and directly depends on feasible paths. When the size of the network (the number of links and nodes in a representative graph) increases, allowing us to explore more paths, the number of feasible paths also increases, and the OD demand matrix may grow accordingly, leading to a non-linear increase in computation time (Patriksson, 2015). Because changing the input (demand and graph) results in a new problem to be solved, it requires re-calculation. Therefore, there is a critical need to find another way to capture the complex nature of traffic flow and demand in large networks in order to use the already-solved problem to estimate the solution to the new problem.
@@
Mathematically, the UE, is typically calculated using optimization methods such as the gradient method or fixed-point algorithm (Tan et al., 2020; Mehrabipour and Hajbabaie, 2021; Liu et al., 2023), e.g., Frank-Wolfe algorithm (Fukushima, 1984). While these optimization-based approaches provide near-optimal solutions, they require solving the entire network problem from scratch for each ‘what-if’ scenario, making them computationally expensive and impractical for real-time or large-scale applications. However, with the presence of data-driven approaches and the application of machine learning, an alternative solution to this problem could be prediction instead of optimization. In other words, we can employ supervised learning to learn the optimal solution and predict it given a new demand or a new graph. Recently, a few studies have tried to estimate the link flow solution of UE using machine learning (Fan et al., 2022; Rahman and Hasan, 2023; Liu and Meidani, 2024; Liu et al., 2025a; Liu and Meidani, 2025; Hu and Xie, 2025) with graph neural networks (GNNs) emerging as a particularly popular approach for utilising network topology to capture spatial dependencies. For example, Rahman and Hasan (2023) and Liu et al. (2025a) use graph convolutional neural networks (GCN) to forecast traffic flow patterns, while Liu and Meidani (2024); Hu and Xie (2025) employed graph attention networks (GATs) to accelerate assignment across large-scale planning scenarios. While promising, GNNs are constrained by their reliance on local message passing, which limits their ability to model long-range dependencies. Deeper stacking can address this, but at the cost of higher computation and over-smoothing of features.
@@
Other approaches have pursued surrogate formulations of the traffic assignment problem. For instance, Zhang and Qian (2025) introduced a low-rank approximation for path-based network models, generating hypothetical path sets based on flows rather than topological structure. Liu et al. (2023) propose an end-to-end learning framework that embeds the traffic equilibrium problem as a Variational Inequality (VI) layer, producing flows through an iterative equilibrium procedure. Yet these methods remain tied to surrogate calibration or traditional iterative equilibrium procedures. While Deep Neural Networks (DNNs) have proven effective in domains such as computer vision and NLP Wang et al. (2022), and are increasingly applied to transportation forecasting, the use of sequence-based architectures, remains largely unexplored in TAP (Wen et al., 2023; Xue et al., 2025; Jia et al., 2025). Recent studies have applied attention mechanisms to enhance sequential modeling in traffic forecasting, such as hierarchical attention-based LSTM frameworks for network-level traffic state prediction Zhang (2024), demonstrating the ability of attention to capture multi-scale temporal dependencies. Building on this, Transformer-based architectures have been explored in related transportation applications involving spatiotemporal time series prediction, including trajectory and motion forecasting Sun et al. (2024) and traffic flow prediction Ji et al. (2025), demonstrating their effectiveness in modeling long-range dependencies and global patterns in traffic data. Despite these advancements, the application of Transformers to the TAP remains largely unexplored.
@@
To address these limitations, we introduce a Transformer-based framework for TAP that employs global self-attention, enabling each element of the input sequence to dynamically weigh its relevance to all others and capture complex, long-range dependencies across the network. The attention mechanism allows the model to selectively prioritize OD pairs with nonzero demand or links with critical features while maintaining global context, enhancing learning efficiency. In contrast, prior TAP models using GAT-based encoder-decoder architectures restrict attention to local neighborhoods and decode flows through MLP layers, limiting their ability to model global interactions (Hu and Xie, 2025). The proposed model offers both computational efficiency and robustness, providing a scalable alternative to traditional iterative equilibrium solvers. To the best of our knowledge, this represents the first TAP framework to employ a Transformer-based sequence learning architecture integrating both encoder and decoder global attention.
Travelers make decisions at the path level, and equilibrium arises from their collective route-choice interactions. Most existing approaches emphasize link-level predictions, leaving path-level demand propagation underexplored, despite its importance for capturing traveler behavior. In contrast, the proposed path-based Transformer learns the equilibrium directly in path-flow space, aligning with the game-theoretic interpretation of TAP. In doing so, the model reformulates TAP as a traveller decision-making process rather than only a traffic prediction problem. By providing detailed path-level flows, it captures the interactions among OD pairs that give rise to equilibrium and provides an interpretable mapping between demand and network conditions. This reformulation addresses two persistent challenges in current practice, the computational cost of iterative solvers and the lack of travelers choice representation in link-based models. Trained on feasible equilibrium data, the model inherently preserves flow conservation and capacity constraints while supporting scenario-based evaluations such as network or demand perturbations, enabling generalization across unseen conditions. This formulation also establishes a foundation for route-choice guidance or decision-support applications in TAP.
@@
To summerize, the major contributions of this work are as follows: (i) A Transformer-based model is proposed for directly learning and predicting UE path flow distributions in both single- and multi-class traffic networks; (ii) The model incorporates both network topology and detailed OD demand information, allowing it to generalize to unseen demand patterns and network changes; (iii) Our approach eliminates the computationally intensive optimization step, providing solutions orders of magnitude faster than traditional methods; and (iv) The model supports robust ’‘what-if" scenario analysis under extreme perturbations, where link-removal and link addition based topology changes, missing-link observation cases, as well as multi-class and large-scale network applications particularly in terms of OD demand are considered jointly with incomplete OD demand.
@@ 2 Related literature
Beginning with UE principle of Wardrop (1952) and the convex optimisation formulation of Beckman et al. (1956), these approaches laid the foundation of the TAP. To reduce the need for full path enumeration, some analytical approaches developed link-based algorithms, which directly solve for link flows (Fukushima, 1984), as well as column and bush-based methods, which aim to find only promising subsets of paths to improve efficiency (Ford and Fulkerson, 1958; Bar-Gera, 2002; Dial, 2006). Despite these advances, TAP remains computationally intensive for large networks, where evaluating thousands of ‘what-if’ scenarios can take days. This motivates the shift toward supervised machine learning models trained on solver-generated solutions, offering significantly faster inference while retaining the accuracy of the underlying optimization framework.
@@ 2.2 Deep Learning (DL) models
Besides the calculation of UE by optimization approaches, there are multiple deep learning approaches for traffic estimation, which can be related to estimating the state of the network. RNNs have been widely used to model temporal dependencies in traffic flow data for prediction. Studies like those by Fernandes et al. (2020) and Wu et al. (2023) have shown the ability of Long Short Term Memory (LSTM) networks, a type of RNN, to capture long-term and short-term dependencies in traffic patterns. More recently, hierarchical attention-based extensions of LSTM models have been proposed for network-level traffic state prediction Zhang (2024), effectively capturing multi-scale temporal dependencies and improving the forecasting of congestion patterns. However RNNs are recursive and cannot parallelize all sequences in parallel, and they are weak for variable-length data (Ma et al., 2015; Bogaerts et al., 2020). Fan et al. (2022) had built a CNN-based model to estimate link flow from incomplete OD data. However, this model was applied on a small-sized dataset of 1000 data points and it does not address the estimation tasks of path flow distribution.
@@ 3 Comparison Between the Proposed Model and Other Neural Network Models
Model Type (Reference) Model Output Baseline OD Data Perturbations Conservation Handling
@@
As discussed in previous sections, the application of neural networks to TAP remains relatively underexplored. Table 1 compares our approach with other state-of-the-art neural network based traffic assignment models in terms of model outputs, handling of perturbations, and flow conservation strategies. Beyond introducing a new architecture, this study empirically demonstrates that a Transformer-based sequence learning formulation can reliably replicate User Equilibrium conditions, preserve OD-level conservation, and sustain performance under simultaneous demand and topology perturbations.
@@
While link or path-level flow conservation ensures locally feasible flows, it does not necessarily guarantee that OD demand is satisfied. To address this, Hu and Xie (2025) applied a post-processing step to adjust path flows and enforce OD demand consistency. In contrast, our path-based Transformer model, trained on ground-truth path flows generated under equilibrium conditions, predicts flows that inherently sum to each OD pair’s demand, ensuring OD-level conservation without any additional adjustment. Because OD conservation is implicit and enforced through formulation in Equation 3, no iterative enforcement or adjustment is needed at inference, making predictions potentially faster.
@@ 4.1 Problem formulation
As mentioned before, the goal is to estimate the optimal solution of path flow distribution from given OD demands and the graph instead of solving it. In mathematical terms, we consider a network represented as a graph $G=(V,E)$, where $V$ stands for the set of nodes with size $N$, and $E$ for the set of links. Let $R$ represent the set of OD pairs. For each OD pair, denoted by $r\in R$, where $r=(v_{1},v_{2}),\forall v_{1},v_{2}\in V$, vehicles of class $z\in Z$ can travel via several finite and nonempty feasible paths of path set $\mathbf{P}_{r}$. Let $n$ denote the number of classes in the network. Each set $\mathbf{P}_{r}\in\mathbf{P}^{k}$ encompasses $k$ feasible paths for each OD pair $r$. Furthermore, let $\mathbf{X}$ represents OD demand matrix, $x^{z}_{r}\in\mathbf{X}$ represents the demand of OD pair $r$ traveled by class $z$, and $f^{r,z}_{\mathbf{P}_{r}}\in\mathbf{F}$ represents paths flow distributions of OD pair $r$ on path set $\mathbf{P}_{r}$, with class $z$. Each set $f^{r,z}_{\mathbf{P}_{r}}$ encompasses $k$ flow values for every OD pair $r$.
@@
The ultimate goal is to find $\mathbf{F}^{*}$, the optimal path flow distribution for UE condition defined by Wardrop (1952). Each set of path flow distribution ${f^{r,z}_{\mathbf{P}^{r}}}^{*}\in\mathbf{F}^{*},\forall r,\forall z$ is the optimal solution for OD pair $r$ on path set $p_{r}$ with class $z$. Equation 1 shows that for all classes $z\in Z$, the paths connecting any OD pair $r$ can be categorized into two groups: those that carry flow, where the travel time $c^{r,z}_{p}$ matches the minimum OD travel time $u^{z}_{r}$, and those that do not carry flow, where the travel time is greater than or equal to the minimum OD travel time. Equation 2 ensures that $u^{z}_{r}$ is the minumum path cost. Equation 3 ensures that all the demand is served. Equation 4 ensures the positivity of variables. However, this UE solution is only applicable when all drivers have perfect knowledge of the network state. In reality, it is not the case, so normally there will be a delay in the network. In the literature, this imperfection of knowledge is represented by stochastic terms that lead to other equilibrium terms such as Stochastic User Equilibrium (SUE) or Boundedly Rational UE (BRUE) (Ameli et al., 2020b).
@@ 5 Numerical Experiments
Two numerical experiments are conducted to evaluate the accuracy, efficiency, and generalization capability of the proposed framework. In Section 5.3, We first experiment with a synthetic Manhattan-like network to assess the basic effectiveness of the model. We evaluate the model’s performance under conditions of incomplete OD demand. The second experiment in Section 5.4 is on urban transportation networks (Sioux Falls and Eastern-Massachusetts network) with several ‘what-if’ scenarios to examine robustness against various realistic network perturbation scenarios. The implementation details and datasets used in these experiments are available at https://github.com/sulthanashams/DL_TAP/tree/main. The details of the experiments will be explained in the following sections.